OpenStudy (anonymous):

What is the equation of the line that passes through the point (1, 6) and has a slope of 3? y = 3x + 5 y = 3x + 3 y = 3x − 17 y = 3x − 5

OpenStudy (anonymous):

solve using formula (y-y1)=m(x-x1)

Parth (parthkohli):

Point-slope *whistles* :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I love point slope form!

Parth (parthkohli):

Yes, can you do it now? =)

OpenStudy (karatechopper):

Dude these problems are so much fun!!!!! Point slope- y-y1=m (x-x1) Now to plug in values!! y-6= 3(x-1) Solve!! y-6=3x-3 y-6+6=3x-3+6 y=3x+3

Parth (parthkohli):

(Careful, that's an answer...) :)

OpenStudy (anonymous):

what kind of sick twisted fun do you like lol, thank you... I have so much trouble with mat and history... I cant remember anything or sit down for long periods of time!

Parth (parthkohli):

@vcbvc: I'm hyperactive as well :)